OpenStudy (anonymous):

A 2.50 kg ball moving at 7.50 m/s is caught by a 70.0 kg man while the man is standing on ice. How fast will the man / ball combination be moving after the ball is caught by the man? ***How would I do this problem? Thank you!

ganeshie8 (ganeshie8):

you may use conservation of momentum \[\large \rm{m_1v_1 = m_2 v_2}\]

ganeshie8 (ganeshie8):

\[\large \rm{2.5\times 7.5 = 70v_2}\] solve \(\rm{v_2}\)

OpenStudy (anonymous):

I can help u

OpenStudy (phi):

ball caught by a man so mass2 is the sum of the man and ball. in other words \[ m_1 v_1 = (m_1 + m_2) v_2 \\ 2.5 \ kg \cdot 7.5 \ m/s= (2.5 + 70) \ kg \cdot v_2 \]

OpenStudy (anonymous):

okay, thank you all!! so would it be v2=0.267857? or is it v2=0.25862? not sure! and i round it to v2=0.26? and so this is how fast? 0.26 m/s ?